OpenStudy (anonymous):

17! / 4! * 13! = A. .0238 B. 23,800 C. 2,380 D. Some number less than 1

OpenStudy (across):

\[\frac{17!}{4!\cdot13!}=\frac{17\cdot16\cdot15\cdot14\cdot\not{13!}}{4!\cdot\not{13!}}=\frac{17\cdot16\cdot15\cdot14}{4!}\]

OpenStudy (anonymous):

17 * 16 * 15 * 14 / 4 * 3 * 2 * 1 17 * 2 * 5 * 14 = 2380

OpenStudy (anonymous):

: C

OpenStudy (turingtest):

@meow18 do you want some more explanation ?